OpenStudy (anonymous):

simplify: 30p^4q-45p^3q^2+15p^2q^3/5p^2q a. 6/p^2-9p/q+3q^2 b. 6p^2-9pq+3q^2 c. 6p^2-9p/q+3q/p d. 25p^2-40pq+10q^2

OpenStudy (anonymous):

is a easy problem

OpenStudy (anonymous):

recall that multiplying is is(x^2)*(x*2)= x^2+2 =x^4 and division is the opposite of mult x^2/x^2= 1

OpenStudy (anonymous):

soo now what?

OpenStudy (anonymous):

\[30p^{4q}-45p^{3q^{2}}++15p^{2q^{3}} / 5p^{2q}\]

OpenStudy (anonymous):

yes i know but i dnt know which answer it is

OpenStudy (anonymous):

if u divide them is 6-9+3 and subtract the power is p^2q+p^q+q^2

OpenStudy (anonymous):

i cant decide if its a or b?

OpenStudy (anonymous):

wait is it all divided by 5p^2q? or jus the last term?

OpenStudy (anonymous):

doesnt make any sense dive p/q

OpenStudy (anonymous):

yes is it \[5p^{2q}\]

OpenStudy (anonymous):

i think its b, im starting to get lost

OpenStudy (anonymous):

45p^3q^2/5p^2q= 9 \[p^{3q^2}\]

OpenStudy (anonymous):

i think its b but im not 100% sure just because how you have the equation set up

OpenStudy (anonymous):